What is the sum of the series 0.5, 2.5, 12.5, to 7th terms?
Answer by stanbon(75887)

What is the sum of the series 0.5, 2.5, 12.5, to 7th terms?
Geometric Series with a = 0.5 and r = 5
----
S(7) = 0.5[(5^6-1)/(5-1)] = 0.5[15624/4] = 3906/2 = 1953
----------
Cheers,
Stan H.
----------